# Lambda-Skip Connections: the architectural component that prevents rank collapse

### v1

@@ A.11 Proof of Counterexample for Selective SSMs
A.12 Additional Experiments
A.12.1 Comparison of Rank Collapse for the S4 Architecture with and without skip connection and LayerNorm
@@
Figure 4 shows that for S4D, even when removing the skip connection and the LayerNorm from the model, rank collapse is not observed. This can be explained by the fact that the condition on the Frobenius norm of the weight matrix being smaller than 1 to guarantee rank collapse is quite stringent in practice and hence it is not met by the model. Additionally, note that while for Transformers and for Mamba-2 the observation of rank collapse depended (at least theoretically) on both the values of the weight matrices and the inputs values, for standard SSMs (e.g. S4) rank collapse only depends on the weight matrix. This renders the SSM model more ”robust” to rank collapse, in the sense that to observe rank collapse one needs to find a precise parametrization of the model and cannot simply find an ”adversarial” input sequence causing rank collapse (which instead could be done for every parametrization of Transformers and selective SSMs).
A.13 Experimental Details

### v3

@@ A.11 Proof of Counterexample for Selective SSMs
A.12 Proof of the Tightness of the Lower Bound in Theorem 4.1
See 4.3.2
Proof.
We again consider [Sys-2] from Proposition 4.3.1 with $\alpha_{0}=\beta_{0}=\frac{1}{2}$. First, we calculate the value of the rank collapse metric at every layer. Following the proof of Proposition 4.3.1, where we saw that $Y^{(k)}=\begin{pmatrix}1&0\\ \frac{(2+\lambda)^{k}\alpha_{0}}{\sqrt{(2+\lambda)^{2k}\alpha_{0}^{2}+(1+\lambda)^{2k}\beta_{0}^{2}}}&\frac{(1+\lambda)^{k}\beta_{0}}{\sqrt{(2+\lambda)^{2k}\alpha_{0}^{2}+(1+\lambda)^{2k}\beta_{0}^{2}}}\end{pmatrix}$, a simple calculation leads to $\mu(Y^{(k)})=\frac{1}{1+(\frac{2+\lambda}{1+\lambda})^{2k}}$. We are now interested the value of $a$, which can be simply calculated by $a=\frac{\mu(Y^{(k+1)})}{\mu(Y^{(k)})}=\frac{1+(\frac{2+\lambda}{1+\lambda})^{2k}}{1+(\frac{2+\lambda}{1+\lambda})^{2k+2}}$.It is easy to see from the previous expression that $a=O(\min(1,(\frac{1+\lambda}{2+\lambda})^{2}))=O(\min(1,\frac{|\lambda|}{1+|\lambda|})$. In particular, we notice that in the region where the model suffers from rank collapse in the limit of infinite depth (i.e. $\lambda>-\frac{3}{2})$), we have that $a=O(\frac{|\lambda|}{1+|\lambda|})$. Note that this corresponds to choosing $|\lambda|=\Omega(\frac{a}{1-a})$, hence concluding the proof. ∎
A.13 Additional Experiments
A.13.1 Comparison of Rank Collapse for the S4 Architecture with and without skip connection and LayerNorm
@@
Figure 4 shows that for S4D, even when removing the skip connection and the LayerNorm from the model, rank collapse is not observed. This an be explained by the fact that the weight matrices at different layers have a large Frobenius norm. From Theorem A.8, in order to have rank collapse, we must have that the weights matrices have Frobenius norm smaller than 1. Hence, the fact that we do not observe rank collapse does not violate our theory as in the case of weight matrices with large Frobenius norm, the upper bound presented in Theorem A.8 becomes vacuous. Additionally, note that while for Transformers and for Mamba-2 the observation of rank collapse depended (at least theoretically) on both the values of the weight matrices and the inputs values, for standard SSMs (e.g. S4) rank collapse only depends on the weight matrix. This renders the SSM model more ”robust” to rank collapse, in the sense that to observe rank collapse one needs to find a precise parametrization of the model and cannot simply find an ”adversarial” input sequence causing rank collapse (which instead could be done for every parametrization of Transformers and selective SSMs).
A.13.2 Comparison of Performance on Image LRA for Different values of Skip Strength
In the following, we test and report the performance for different values of $\lambda$ for Transformers for the Image LRA task. In particular, we report a table with the performance of Transformers (with 8 layers) trained from scratch on the Image LRA dataset for different values of $\lambda$, namely -2, -1, 0, 1, 2. From the results, we can make the following conclusions: we can clearly see that rank collapse is happening for the case where $\lambda=0$, proving the importance of having skip connections in the architecture. Indeed, the model for this value of $\lambda$ performs random guessing, meaning that it is not able to build useful representations of the inputs due to rank collapse. For the other cases, the performance is comparable across different values of $\lambda$.
Model $\lambda=-2$ $\lambda=-1$ $\lambda=0$ $\lambda=1$ $\lambda=2$
Transformers 38.61 35.89 10.00 40.22 38.90
Figure 5: Performance metrics of Transformer and Mamba-2 for different values of $\lambda$.
A.14 Experimental Details